OpenStudy (sugarapples):

g(x)=x+4 y=x+4 may I a please get help

OpenStudy (anonymous):

hi... okay but what is the Q

OpenStudy (sugarapples):

g(x)=x+4

OpenStudy (anonymous):

okay... g(x)=x+4 then !

OpenStudy (solomonzelman):

g(x)=x+4 and y=x+4 is the same thing

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

y=x+4 to find the inverse, switch x and y, and subtract 4 from both sides.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y=x-4

OpenStudy (anonymous):

okay i have to go now